Please help me graph these two equations on a number line and write in interval notation….
1) |4-3x| < or equal to 13
2) |2x-5|-5 > 20


Can you please also explain the graphs on these two equations. I'm really trying to understand how to graph these. Some of them look really weird, so it seems incorrect. Please help.
Thank you So much!

Answer by stanbon(75887)

Please help me graph these two equations on a number line and write in interval notation….
1) |4-3x| <= 13
-13 <= 4-3x <= 13
-17 <= -3x <=9
-3 <= x <= 17/3
Draw a number line
Plot x = -3 and x = 17/3 = 5 2/3
Put a solid circle at each of those values
Draw a line between the solid circles to see the total graph
Interval:: [-3,17/3]
==================================
2) |2x-5|-5 > 20
|2x-5| > 25
2x-5 > 25 or 2x-5 < -25
2x > 30 or 2x < -20
x < 10 or x > 15
Draw a number line; plot x = 10 and x = 15
Draw an open circle at x = 10 and at x = 15
Draw an arrow to the left from x = 10
Draw an arrow to the right from x = 15
Interval:: (-oo,10)U(15,oo)
===================
--------------------
Cheers,
Stan H.
----------------